OpenStudy (calculusxy):

MEDAL!!! Write a system of equations with the solution set of (3, 4). Make the linear equations have a positive slope.

OpenStudy (patrickjbraddr):

one second

OpenStudy (patrickjbraddr):

crap I'm the wrong person ask @Vocaloid

OpenStudy (calculusxy):

@mathstudent55

OpenStudy (calculusxy):

@Vocaloid

Vocaloid (vocaloid):

it doesn't say how many or what kind, so let's keep it simple

Vocaloid (vocaloid):

a linear equation with a positive slope will have the form -__x + y = ___ where two positive numbers go into the blanks

Vocaloid (vocaloid):

so you just wanna write two equations that work when you plug in x = 3 and y = 4

Vocaloid (vocaloid):

ex: -2x + y = -2 let me know if you need help coming up with another one

Vocaloid (vocaloid):

*also i made an error before, only the first blank has to be positive

OpenStudy (calculusxy):

@Vocaloid I need help coming up with the other one :)

Vocaloid (vocaloid):

-__ * 3 + 4 = ___ pick a number for the first blank and see what the second blank should be equal to